OpenStudy (anonymous):

will give medal Simplify the expression. (10 + 2) – 2 1 5 A. 146 B. 75 C. 125 D. 96

OpenStudy (jadeishere):

distribute 2 to 10 and 1 so is it (10 + 1) squared or multiplied by two?

OpenStudy (jadeishere):

if it's (10 + 1) ^2 - 52 then it'd be (10 + 1)(10 + 1) - 52 100 + 2 - 52

OpenStudy (anonymous):

it is (10+1)^2 - 5 ^2

OpenStudy (jadeishere):

Ohhhh, makes more sense!! (10 + 1)^2 - 5^2 then do PEMDAS 11^2 = 121 5^2 = 25 121 - 25 = 96

OpenStudy (anonymous):

thx

OpenStudy (jadeishere):

Yep